OpenStudy (helplilly12):

Use a related equation that has the variable on one side. Then simplify the other side. D/6=2 A.d=12 B.d=8 C.d=4 D.d=1/3 Please...help im tired;-; T^T

OpenStudy (mathgeek289):

The Answer Is A, Because: D/6 = 2 Multiply Both Sides By 6. (D/6) * 6 = 2 * 6 (D/6) * 6 = D 2 * 6 = 12 D = 12

OpenStudy (helplilly12):

thank you

OpenStudy (mathgeek289):

Medal Or Best Response Is Very Appreciated!!